This is Imaginary Solutions:
(y-2/3)^2=4/9

Answer by richwmiller(17219)

Huh?
(y-2/3)^2=4/9
y-2/3=+-2/3
y=2/3-2/3=0
y=4/3=1 1/3
Nothing imaginary here